OpenStudy (anonymous):

Can anyone help? Solve |4x-11|<5 Showing all work. In complete sentences describe the graph in words.

OpenStudy (ivettef365):

you need to solve is in 2 ways, as follows: 4x - 11 < 5 and 4x - 11 > -5 (solve for x)

OpenStudy (anonymous):

Thanks! I wasnt really sure if the signs would be switched or not.

OpenStudy (ivettef365):

yw :)

OpenStudy (anonymous):

I got x = 4 and x= 6/4 as my final answers, is that correct?

OpenStudy (ivettef365):

yes, and you can simplify the 6/4 = 3/2

OpenStudy (anonymous):

Oh, completely missed that bit, thank you. The graph will be positive, right?